# Resolving the Exploitation-Exploration Dilemma in Evolutionary Algorithms: A Novel Human-Centered Framework

### v1

@@ 2.3 Some Key Insights
    Efficient search navigation in a GA fundamentally relies on achieving a balanced exploration-exploitation dynamic.
@@
    Exploitation and exploration in a GA can be addressed on two distinct levels: the semantic level and the intensity level. The semantic level focuses on the descriptive content of the mechanisms or strategies used to achieve exploration and exploitation (the "how"). In contrast, the intensity level determines the degree to which these mechanisms are applied (the "how much"). Together, these levels shape the algorithm’s ability to balance searching diverse areas of the solution space with refining promising solutions.
@@
    The traditional internal operators (selection, crossover, mutation, replacement) and parameters (e.g., population size, crossover probability, mutation rate) collectively define the GA’s potential for exploration and exploitation.
@@
    Internal exploration and exploitation processes in a GA, driven by traditional operators and parameters, are inherently conflicting. Optimizing one often diminishes the other
@@
    The structure of GA tells us that the algorithm operates and moves within the representation set $X$. The cardinality of $X$ is $|X|=K^{L}$ where $K$ is the number of representative elements and $L$ is the length of the string.
@@
    The structure of GA is abstract enough to give the user flexibility in defining their own operators and parameters, enabling customization and adaptation to specific problem domains.
  7. 7.
    Parameter tuning in optimization algorithms is a crucial task, as it significantly impacts algorithm behavior, convergence, and solution quality. It remains an open question in research to continuously discover new parameters tailored to specific problem domains, leading to enhanced optimization outcomes and novel insights.
@@ 3 Managing the Exploration-Exploitation Dilemma in EAs: The Current Approach
Building upon the seventh paragraph in the introduction section, this section analyzes the limitations of traditional approach for enhancing exploration without compromising exploitation in GAs. Although GAs are used as a representative case, these limitations are applicable to the broader class of EAs.
To illustrate these limitations, let us consider the idealized case in which no point in the search space is revisited throughout a single execution of the algorithm.
Increasing the exploration potential of a GA requires fulfilling two necessary conditions: (1) increasing the size of the representation set and (2) the size of the population. As mentioned in section (2), the size of the representation set within which the GA operates is $K^{L}$, where $K$ is the base of the representation system (e.g., binary (2), ternary (3), decimal (9), etc.) and $L$ is the length the representative string of each potential solution. So, increasing the size of the representation set, for some fixed representation system, implies increasing the value of $L$.
Assume that the algorithm explores $N$ solutions per iteration, and given the assumption that no point in the search will be visited twice (which is in practice, not guaranteed to happen) this would imply the need for $m$ number of iterations where $m$ satisfies $\left|m-\frac{K^{L}}{N}\right|\leq\epsilon$ and $\epsilon=\min\left\{y\in\mathbb{Z^{+}}\mid\left|m-\frac{K^{L}}{N}\right|\leq y\right\}$ for the algorithm to explore the whole representation set.
@@
It can be inferred that with the addressed approach, enhancing the exploration potential not only entails satisfying conditions (1) and (2) but also demands meticulous tuning of internal GA operators (e.g. crossover + mutation) to facilitate the exploration of new search points as much as possible and minimize the risk of getting trapped in a narrow and suboptimal region of the solution space but without risking it behaving like a random algorithm.
Moreover, several challenges arise immediately:
@@
  • The second challenge stems from the inherent unintelligibility of the interdependencies between internal GA parameters and operators which further complicates the understanding and prediction of their combined effects on the exploitation-exploration trade-off thereby, making their tuning process difficult.
  • The third challenge is that even if conditions (1) and (2) are met, and a selection of good GA operators was employed, there are no guarantees that the algorithm will reach promising solutions. It is possible that some promising niches may not be included within the bounds of the workspace  [^4] The pre-image of the representation set under the DECODE map, if we are following the DECODE-ENCODE strategy, or the domain of the ENCODE map, if we are following the ENCODE-DECODE strategy of the algorithm, leading to the risk of overlooking optimal solutions.
The limitations identified in the current approach are not exclusive to GAs but are inherent to the broader class of EAs. This is because all EAs share a common foundation: they are population-based methods that encode candidate solutions using structured representation systems and rely on stochastic operators to guide the search process.
Given these challenges, it is clear that an alternative approach is needed to address the exploitation-exploration dynamics more effectively.
@@ 4 A Novel Human-Centered Two-phase Search Framework (HCTPS)
In this section, we propose a human-centered framework, designed to enhance the exploration of the search space while preserving the exploitation capability of an EA, ensuring it does not devolve into random behavior. This framework is structured as a two-phase search process, consisting of:
@@
    Global Search Phase: In this phase, a broad exploration of the entire search cube $I$ is conducted. In this, the algorithm’s internal exploration-exploitation potential is distributed across the whole search space. However, this phase alone may overlook certain promising regions or niches within $I$, necessitating a subsequent localized search to address this limitation.
@@
    Local Search Phase: In this phase, a sequence of selected sub-cubes $\{I_{i}\}$ from the original search cube $I$ (where $i=1,2,\dots,n$) undergoes further intensive sequential exploration. In this, the algorithm focuses its internal exploration-exploitation potential over each sub-cube of the generated sequence.
    The rigorous implementation of the local search phase is achieved through the introduction and utilization of the human-centered search space size control parameter (HSSCP) which serves as a defining mechanism for the selection and exploration of the $n$ specific sub-cubes within $I$.
4.1 The Human-Centered Search Space Size Control Parameter (HSSCP)
The HSSCP is an external configuration variable to an EA that represents a selection of a sequence of sub-cubes $\{I\}_{i=1}^{n}\subset I$ for targeted exploration during the local search phase.
The following steps outline how the decision-maker sets, tunes, and updates the HSSCP:
@@ 4.1 The Human-Centered Search Space Size Control Parameter (HSSCP)
    Rule Specification: Before the local search phase begins, the decision-maker establishes a rule to generate the sequence of sub-cubes $\{I\}_{i=1}^{n}$.
@@
    Sequence Generation: Subsequently, the decision-maker generates the first $n$ terms of this sequence according to the specified rule, where each term corresponds to a localized region of interest that the algorithm will intensively explore.
Throughout this process, the decision-maker retains complete flexibility in determining the number of sub-cubes $n$ to be explored. They also have the autonomy to generate additional terms in the sequence as needed, enabling real-time adjustments based on the optimization problem’s complexity, insights gained during exploration, and individual preferences.
4.2 HCTPS: Pseudo-Algorithm
The presented pseudo-algorithm outlines the general framework of the two-phase search process with HSSCP integration.
The specific implementation details may vary based on the algorithm $g$ used, and the characteristics of the output data, including whether it constitutes a single solution set, a multi-solution set (such as a Pareto set in case of multiobjective optimization), or descriptive statistics measures .. etc.
Algorithm 1 HCTPS Algorithm
@@ 4.2 HCTPS: Pseudo-Algorithm
     Original search cube $I$, HSSCP sequence length $n$, evolutionary algorithm $g$,
     stopping criteria: decision-maker satisfaction $T_{\text{satisfactory}}$
@@
     The best results found during the entire exploration process to $T_{\text{satisfactory}}$
@@
 a) Run $g$ across the entire search cube $I$ to perform a broad exploration experiment.
 b) Record the output results.
Step 2: Local Search Phase with HSSCP:
 a) Specify the rule to generate a sequence of sub-cubes $\{I\}_{i=1}^{n}$ within $I$,
 and generate the first $n$ terms of the sequence, where each term represents a localized region of interest
 b) Run $g$ through each sub-cube to conduct an intensive exploration.
 c) Record the output results of each $i$ exploration experiment.
Step 3: Additional Sub-Cubes Generation:
while the stopping criteria $T_{\text{satisfactory}}$ is not met do
      a) Generate $n_{\text{additional}}$ new terms in the sequence and go to step 2c.
@@
The value of $n_{\text{additional}}$ represents the number of additional terms in the sequence that can be generated after the $n$ exploration experiments are carried out in real-time. This allows the decision-maker to dynamically adjust the exploration based on insights gained, problem complexities, and satisfaction level.
@@ 4.3 Visualizing the HCTPS with a Helicopter Analogy
This dynamic approach intelligently compensates for the camera’s fixed zoom capability through the human-centered adjustment of the helicopter’s position, allowing exploration of the landscape from various positions and angles. In this scenario, the camera’s fixed zoom capability is analogous to the algorithm’s internal exploitation-exploration fixed potential. The 3000 feet view is analogous to the initial search phase (the global phase) in our proposed approach and this is where traditional EAs stop at thereby, only able to explore the search space from a fixed perspective.
@@
5 GA-integrated HCTPS: Experimental Results
To evaluate the effectiveness of the proposed framework, we compare the performance of the standalone GA with the GA integrated into the HCTPS framework described by algorithm (1) where $g$ is the canonical GA on fourteen challenging black-box, unconstrained optimization problems from ([16], [17], [18], [19]), focusing exclusively on their 30-dimensional versions.
These problems exhibit varying levels of difficulty and topological complexity. For instance, functions like F1 (Bent Cigar) are relatively simple and serve as a baseline for performance comparison, while others, such as F10 (Rastrigin) and F11 (Griewank), present significant challenges due to their multimodal nature and the presence of numerous local optima. Moreover, the set contains functions that are specifically designed to be computationally expensive such as F14 (Ackley), thereby simulating real-world scenarios where evaluations of the objective function are costly ([20]), and providing a realistic framework for assessing the effectiveness of various exploration and exploitation EA strategies to navigate complex landscapes ([21]. Detailed discussions of these functions’ properties are provided in ([22], [16]).
Our GA operates with a maximum limit of exact objective function evaluations per run, defined as $\textit{feval}=50\times n$, where $n$ is the dimensionality of the problem. For the 30-dimensional case, $\textit{feval}=1500$. The search cube for all problems is $I=[-100,100]^{30}$, which includes the global minimum for each test problem as well. The GA employs binary encoding, two-point crossover, and tournament selection. Statistical results are derived from 20 independent runs.

### v2

@@ 3.2 Some Key Insights
    Efficient search navigation and convergence in an EA fundamentally relies on achieving a properly balanced exploration-exploitation dynamic relative to the problem environment.
@@
    The traditional internal operators (e.g., selection, crossover, mutation, replacement) and parameters (e.g., population size, crossover probability, mutation rate) collectively define the EA’s inherent potential for exploration and exploitation.
@@
    The exploration and exploitation processes driven by such traditional operators and parameters, are inherently conflicting. Optimizing one process often diminishes the other.
@@
    The structure of an EA tells us that the algorithm operates and moves within the representation set $X$. The cardinality of $X$ is $|X|=K^{L}$ where $K$ is the number of representative elements and $L$ is the length of the string.
@@
    The structure of EAs is abstract enough to give the user flexibility in defining their own operators and parameters, enabling customization and adaptation to specific problem domains.
@@
    It remains an active area of research to explore and identify new parameters that influences the optimization performance of algorithms in noticeable magnitudes with reasonable tuning requirements.
@@ 4 Managing the Exploration-Exploitation Dilemma in EAs: The Current Approach
As highlighted in the introduction section, the current approach predominantly focus on adjusting traditional EA parameters (e.g., evolutionary operators and population size). While these tuning strategies can offer incremental gains under certain conditions, they do not amount to a systematic resolution of the exploration-exploitation dilemma—largely due to the unintelligible interdependencies among these parameters and the resulting inability to decouple exploration from exploitation.
Here, we provide a more detailed analysis and offer a more concrete exploration of why and how the current approach’s limitation persist. To illustrate them clearly, let us consider the case of maximizing exploration in an idealized setting in which the algorithm never revisits any point in the search space during a single execution:
As discussed in sections (2 & 3):
  1. 1.
    EAs operate within the representation set $X$, which defines the space of possible encoded solutions, and where the whole movement of the algorithm occurs.
  2. 2.
    The set $X$ is in one-to-one correspondence with a proper subset of the whole search cube $I$ under the DECODE function. Formally, if $f$ is the DECODE function, then $X$ is in one-to-one correspondence with $f(X)$ where $f:X\to f(X)\subset I$.
[…]
Assume that the algorithm evaluates $N$ solutions per iteration, and given the assumption that no point in the search will be revisited (which is in practice, not guaranteed to happen) this would imply the need for $m$ number of iterations where $m$ satisfies $\left|m-\frac{K^{L}}{N}\right|\leq\epsilon$ and $\epsilon=\min\left\{y\in\mathbb{Z^{+}}\mid\left|m-\frac{K^{L}}{N}\right|\leq y\right\}$ for the algorithm to explore the whole representation set.
@@
It can be inferred that with the addressed approach, enhancing the exploration potential not only entails satisfying conditions (1) and (2) but also demands meticulous tuning of internal EA operators (e.g. crossover + mutation) to facilitate the exploration of new search points as much as possible and minimize the risk of getting trapped in a narrow and sub-optimal region of the solution space but without risking it behaving like a random algorithm.
In this, several challenges arise immediately:
@@
  • The second challenge stems from the inherent unintelligibility of the interdependencies between internal EA parameters and operators which further complicates the understanding and prediction of their combined effects on the exploitation-exploration trade-off thereby, making their tuning process difficult.
  • The third challenge is that even if conditions (1) and (2) are met, and a selection of good EA operators was employed, there are no guarantees that the algorithm will reach promising solutions. It is possible that some promising niches may not be included within the bounds of $f(X)$, leading to the risk of overlooking optimal solutions.
Take away message: By considering idealized scenario in which no point in the search space is revisited, one sees why simply scaling up population size or adjusting operator parameters still fails to yield a principled mechanism for exploration—especially as it omits any notion of how to preserve exploitation performance.
Given these limitations, we adopt a less-traveled path: rather than viewing exploration and exploitation as irreconcilable forces or favoring one at the expense of the other, we attempt to reconcile them under a new, unifying framework.
@@ 5 A Novel Human-Centered Two-phase Search Framework (HCTPS)
5.1 Framework Definition and Human-Centered Role
Rather than increasing the size of $X$ and $N$ or direct all efforts to fine-tuning traditional EA parameters (such as mutation and crossover in GAs) to enhance search exploration, we propose a human-centered search framework. This framework dynamically relocates the representation set’s co-domain under $f$ within the search space (i.e., it adaptively shifts the mapping $f$ to sufficiently distinct subsets of $I$). This ensures an increased exploration of new regions within $I$ without compromising the algorithm’s ability to exploit promising solutions while also eliminating the inefficiencies of the traditional approach.
In subsection (2.2), we formally defined a parameter as a fixed quantity that influences the behavior of a system without being an explicit variable in execution. We further categorized parameters based on how they are tuned, distinguishing between:
  • Human-centered parameters, which are manually adjusted by the user.
  • Mechanically-centered parameters, which are autonomously updated by the algorithm’s internal mechanics.
[…]
By the end of this section, the role of the human user as a meta-parameter will become more concrete.
5.2 Two-Phase Structure of HCTPS
This framework is structured as a two-phase search process, consisting of:
@@ 5.2 Two-Phase Structure of HCTPS
    Global Search Phase: In this phase, a broad exploration of the entire search cube $I$ is conducted. In this, the algorithm’s inherent exploration-exploitation potential is distributed across the whole search cube. However, this phase alone may overlook certain promising regions or niches within $I$, necessitating a subsequent localized search to address this limitation.
@@
    Local Search Phase: In this phase, a sequence of selected sub-cubes $\{I_{i}\}$ from the original search cube $I$ (where $i=1,2,\dots,n$) undergoes further intensive sequential exploration. In this, the algorithm focuses its inherent exploration-exploitation potential over each sub-cube of the generated sequence.
    The rigorous implementation of the local search phase is achieved through the introduction and utilization of the search space size control parameter (SSCP) which serves as a defining mechanism for the selection and exploration of the $n$ specific sub-cubes within $I$.
5.2.1 The Search Space Size Control Parameter (SSCP)
The SSCP is an external configuration variable to an EA that represents a selection of a sequence of sub-cubes $\{I\}_{i=1}^{n}\subset I$ for targeted exploration during the local search phase.
The SSCP follows a human-centered tuning strategy. The following steps outline how the user specifies, updates, and refines the SSCP:
@@ 5.2.1 The Search Space Size Control Parameter (SSCP)
    Rule-specification and sequence generation: Before the local search phase begins, the user defines a rule to generate the sequence of sub-cubes $\{I\}_{i=1}^{n}$. Subsequently, the user generates the $n$ terms of this sequence according to the specified rule, where each term corresponds to a localized region of interest that the algorithm will intensively explore.
@@
    Adaptation and expansion: During the local search phase, it is up to the user to expand the sequence by introducing new sub-cubes , or modify the sequence-generative rule dynamically based in response to real-time insights.
5.2.2 HCTPS Procedure
The following structured procedure outlines the general framework of the two-phase search process with SSCP integration.
The specific implementation details vary based on the algorithm $g$ used and the characteristics of the output data, including whether it constitutes a single solution set, a multi-solution set (such as a Pareto set in case of multi-objective optimization), or another form.
Procedure 1 HCTPS Procedure (Structured Steps)
@@ 5.2.2 HCTPS Procedure
     $I$: Global search cube
     $g$: An arbitrary EA
     $k$: Number of iterations per execution of $g$
     $T_{\text{satisfactory}}$: Stopping criterion based on user satisfaction
@@
     $S_{\text{best}}$: The best solution(s) found during the entire search process up to $T_{\text{satisfactory}}$
@@
 a) Run $g$ across the entire search cube $I$ to perform a broad exploration experiment for $k$ number of iterations.
 b) Identify and store the best solution(s) found in this phase as $S_{\text{global}}$.
Step 2: Local Search Phase
 a) Define a sequence-generative rule for selecting sub-cubes $\{I_{i}\}_{i=1}^{n}\subset I$
 b) Generate these $n$ sub-cubes in the sequence according to this rule
 c) For each selected sub-cube $I_{i}$:
  i) Execute $g$ over $I_{i}$ for $k$ iterations to conduct a local search.
  ii) Identify the best solution(s) within $I_{i}$, store as $S_{i}$
Step 3: Additional Sub-Cube Selection
while $T_{\text{satisfactory}}$ is not met do
      a) Generate $n_{\text{additional}}$ new sub-cubes according to either the predefined rule or an updated rule
      b) Execute step 2c for each new sub-cube
@@
Step 4: Output Best Solution(s)
 a) Aggregate results from all exploration searchs in Steps 1 and 2:
  $S_{\text{HCTPS}}=S_{\text{global}}\cup\bigcup_{i=1}^{n+n_{\text{additional}}}S_{i}$
 b) Identify the best solution(s) $S_{\text{best}}$ from $S_{\text{HCTPS}}$
 c) Return $S_{\text{best}}$
The value of $n_{\text{additional}}$ here represents the number of additional terms in the sequence that can be generated after the $n$ exploration experiments are carried out in real-time in step 2. This allows the decision-maker to dynamically adjust the exploration based on insights gained, problem complexities, and satisfaction level.
@@
In the HCTPS framework, the user focuses all of their tuning efforts on one the SSCP.
5.3 Theoretical Guarantee: HCTPS Ensures Increased Exploration
Theorem 1.
Let $I\subset\mathbb{R}^{d}$ be the search cube, and let $X$ be the representation set such that an EA maps $X$ to an effective search region $I^{\prime}\subset I$ via a fixed decoding function $f:X\to I^{\prime}$.
Define two search approaches:
  • Traditional: The EA is applied using a fixed mapping $f$, meaning the search is restricted to a fixed region $I^{\prime}\subset I$.
  • HCTPS: The search space is divided into sub codomains $\{I_{i}\}_{i=1}^{n}$, and $f$ is dynamically relocated to different sub codomains over successive runs.
Then, HCTPS guarantees greater exploration coverage than the traditional one, meaning the probability of covering more of the search space under HCTPS is strictly greater than under the traditional approach.
Proof.
Let $x^{*}$ be an optimal solution (or a high-quality region containing $x^{*}$) in $I$.
In the traditional approach, the probability of reaching $x^{*}$ is given by:
[…]
where $p(x)$ is the probability density function of the search distribution within $I$, which describes the likelihood of sampling different points.
Assume that $p(x)$ is approximately uniform within $I^{\prime}$, which is reasonable under general EA sampling assumptions. Then
$P_{\text{trad}}\approx\frac{|I^{\prime}|}{|I|}P_{\max},$
where:
  • $P_{\max}$ is the probability of discovering $x^{*}$ if the full space $I$ were searched exhaustively with a uniform sampling strategy.
  • $|I|$, $|I^{\prime}|$ represents the volumes (or measures) of $I$ and $I^{\prime}$ respectively.
  • $\frac{|I^{\prime}|}{|I|}$ represents the fraction of the search space that is covered by the EA.
Since $|I^{\prime}|\ll|I|$, it follows that:
$P_{\text{trad}}\ll P_{\max}.$
This means that in the traditional approach, the probability of discovering an optimal region or a solution is fundamentally limited by the fixed and often small search region $I^{\prime}$.
In the local phase of the HCTPS, instead of being confined to a single fixed region $I^{\prime}$, the user selects a sequence of $n$ subregions (subcubes) $\{I_{i}\}_{i=1}^{n}\subset I$. Each execution of the EA occurs within a different $I_{i}$. This modifies the probability of finding $x^{*}$ as follows:
$P_{\text{local}}=1-\prod_{i=1}^{n}(1-P_{I_{i}}),$
where $P_{I_{i}}$ is the probability of $x^{*}$ being located in subcube $I_{i}$.
Since each $I_{i}$ is sufficiently distinct (i.e., covers new unexplored regions), we approximate $P_{I_{i}}$:
$P_{I_{i}}\approx\frac{|I_{i}|}{|I|}P_{\max}.$
And substitute in the earlier equation:
$P_{\text{local}}\approx 1-\prod_{i=1}^{n}\left(1-\frac{|I_{i}|}{|I|}P_{\max}\right).$
This reflects the increasing probability of finding $x^{*}$ as more subcubes are explored.
Comparing $P_{\text{trad}}$ and $P_{\text{local}}$
Since $P_{\text{trad}}\approx\frac{|I^{\prime}|}{|I|}P_{\max}$, we compare it with $P_{\text{local}}$.
For small $\frac{|I_{i}|}{|I|}P_{\max}$, we use the first-order approximation:
$\prod_{i=1}^{n}\left(1-\frac{|I_{i}|}{|I|}P_{\max}\right)\approx 1-\sum_{i=1}^{n}\frac{|I_{i}|}{|I|}P_{\max}.$
Thus, we obtain:
$P_{\text{local}}\approx\sum_{i=1}^{n}\frac{|I_{i}|}{|I|}P_{\max}.$
Since the selected sequence of subcubes covers a total region of volume: $\sum_{i=1}^{n}|I_{i}|>|I^{\prime}|$, we conclude that:
$P_{\text{local}}>P_{\text{trad}}.$
And because $P_{\text{HCTPS}}=P_{\text{trad}}+P_{\text{local}}$, then $P_{\text{HCTPS}}>P_{\text{trad}}$.
Conclusion. Since $P_{\text{HCTPS}}>P_{\text{trad}}$, it follows that HCTPS guarantees superior search coverage and increased exploration compared to the traditional approach. This result holds as long as the selected $\{I_{i}\}$ provides sufficient diversity.
Thus, HCTPS is mathematically guaranteed to enhance exploration while maintaining inherent exploitation driven by the algorithm’s internal mechanics $\square$
@@ 5.4 Visualizing the HCTPS with a Helicopter Analogy
This dynamic approach intelligently compensates for the camera’s fixed zoom capability through the human-centered adjustment of the helicopter’s position, allowing exploration of the landscape from various positions and angles. In this scenario, the camera’s fixed zoom capability is analogous to the algorithm’s inherent exploitation-exploration fixed potential. The 3000 feet view is analogous to the initial search phase (the global phase) in our proposed approach and this is where traditional EAs stop at thereby, only able to explore the search space from a fixed perspective.
@@
6 GA-integrated HCTPS: Experimental Demonstration
The objective of this experimental study is threefold:
  1. 1.
    Illustrate the implementation of the HCTPS: Show how an EA can be adapted to the two-phase HCTPS framework in practice
  2. 2.
    Demonstrate SSCP tuning: Provide an example into how the user configures and refines the SSCP for a particular problem instance
  3. 3.
    Dominant-parameter illustration: Illustrate how the SSCP is the most dominant parameter, in effect, to effective exploration regulation.
Toward goals (1) and (2), we instantiate the procedure (1) using a canonical GA as our base algorithm $g$. We evaluate this GA-integrated HCTPS (HCTPS-GA) on 14 black-box, unconstrained optimization problems collected from (Chen et al., 2014; Yang, 2010; Dieterich & Hartke (2012); Tan, 2016), focusing exclusively on their 30-dimensional versions.
For (3), we compare the proposed framework against the current one (i.e, we compare the performance of the GA integrated into the HCTPS framework (HCTPS-GA) against the standalone GA). See remark (3) in section (5).
These problems exhibit varying levels of difficulty and topological complexity. For instance, functions like F1 (Bent Cigar) are relatively simple and serve as a baseline for performance comparison, while others, such as F10 (Rastrigin) and F11 (Griewank), present significant challenges due to their multimodal nature and the presence of numerous local optima. Moreover, the set contains functions that are specifically designed to be computationally expensive such as F14 (Ackley), thereby simulating real-world scenarios where evaluations of the objective function are costly (Crico et al., 2020), and providing a realistic framework for assessing the effectiveness of various exploration and exploitation EA strategies to navigate complex landscapes (Vaidya et al., 2020). Detailed discussions of these functions’ properties are provided in (Plevris & Solorzano, 2022; Chen et al., 2014).
Our GA operates with a maximum limit of exact objective function evaluations per run, defined as $\textit{feval}=50\times n$, where $n$ is the dimensionality of the problem. For the 30-dimensional case, $\textit{feval}=1500$. The search cube for all problems is $I=[-100,100]^{30}$, which includes the global minimum for each test problem as well. The GA employs binary encoding, the canonical two-point crossover, and binary tournament selection. Statistical results are derived from 20 independent runs.